OpenStudy (anonymous):

2x-y =6, x+2y=3 : solve the system of equations by graphing. Someone help me understand this please

OpenStudy (anonymous):

you need to get y by it self.

OpenStudy (anonymous):

hi can u help me

OpenStudy (anonymous):

@surdawi

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

Yeah i am really lost sorry all!

OpenStudy (anonymous):

use the first equation to isolate y on the left hand side. Once you have that you can plug that value of y into the second function and solve it for x. Once you have the value of x you can go back to the first fuction to compute y

OpenStudy (anonymous):

Although i still dont understand i will try my hardest

OpenStudy (anonymous):

I think it is all the plus and minus signs plus the variables that are getting me the most

OpenStudy (anonymous):

a) 2x-y =6 b) x+2y=3 from a) 2x = 6 +y 2x -6 = y now we can use that in the b) x + 2 * ( y ) = 3 x + 2 * ( 2x - 6 ) = 3 ... continute to solve for x once you know the value of x you can substitute it back into 2x -6 = y and this will give you the value of y

OpenStudy (anonymous):

So x =3 then I am sorry I am so confused and wasting your time thank you for your help